# Uncertainty Quantification for Gradient-based Explanations in Neural Networks

### v1

@@ 5.3 Explanation Uncertainty - Prediction vs Ground Truth Neurons
$\mu$-IG $\sigma$-IG CV-IG $\mu$-GBP $\sigma$-GBP CV-GBP
Refer to caption Deep Ensemble Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-Dropout Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-DropConnect Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: The visualization of the mean ($\mu$), the standard deviation ($\sigma$), and the coefficient of variation ($CV$) heatmaps generated using the combinations of uncertainty estimation methods (Deep Ensemble, MC-Dropout, MC-DropConnect and Flipout) and explanation methods (Guided Backpropagation and Integrated Gradients). The uncertainty estimation methods are organized in individual rows and the explanation methods are aligned along the columns with the first three columns representing the $\mu$, $\sigma$, and the $CV$ heatmaps for IG and the last three columns representing the same for GBP. The image belongs to the FER+ dataset.
5.4 Explanation Uncertainty for Tabular Regression
@@
It can be observed that the mean explanations generated by LIME generally have a lower activation as compared to GBP. This indicates less agreement amongst the constituent explanations and in turn higher uncertainty in the explanation generated by LIME. Furthermore, we find that for LIME, the value of the coefficient of variation is slightly higher for several features, especially in the case of Deep Ensemble and MC-DropConnect. This implies that these explanations are highly uncertain and should be used with caution. This can be attributed to disagreement among the constituent explanations for that particular feature. The same is observed in the case of GBP for a couple of features, (to a smaller degree as compared to LIME), especially for the case of Deep Ensembles.
@@
Figure 11: The explanations for the tabular regression task with the mean and the standard deviation explanation (rows 1-2) and the coefficient of variation explanations (rows 3-4). The first and the third rows represent the outputs of GBP and the second and the fourth rows represent LIME. From left to right, the columns represent as follows: (i) Deep Ensemble (ii) MC-Dropout (iii) MC-DropConnect (iv) Flipout. The input features mapping in the plots is as follows: [1: longitude, 2: latitude, 3: housing median age, 4: total rooms, 5: total bedrooms, 6: population, 7: households, 8: median income].
@@ 6 Conclusions and Future Work
Both uncertainty estimation and saliency explanations have broad impact on society, as machine learning models are increasingly being used in real-world settings, where humans could be harmed, and uncertainty estimation produces better calibrated output uncertainties to detect when a model makes incorrect or ambiguous predictions.
Similarly for saliency explanations, when explanation uncertainty is produced by using an uncertainty estimation method, this would allow human users to obtain additional information in order to decide if the explanation should be trusted or not.
@@
In both cases, output uncertainty and saliency explanations can be misleading, and always additional validation is necessary when used by human users.

### v2

@@ 5.2 Prediction vs Ground Truth Neurons
Explanation for Predicted Class Explanation for Ground Truth Class
$E_{\mu}$ IG $E_{\sigma}$ IG CV IG $E_{\mu}$ GBP $E_{\sigma}$ GBP CV GBP $E_{\mu}$ IG $E_{\sigma}$ IG CV IG $E_{\mu}$ GBP $E_{\sigma}$ GBP CV GBP
Ensemble Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-D Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-DC Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Flipout Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Visualization of the mean ($E_{\mu}$), the standard deviation ($E_{\sigma}$), and the coefficient of variation ($CV$) heatmaps generated using the combinations of uncertainty estimation methods (Deep Ensemble, MC-Dropout, MC-DropConnect and Flipout) and explanation methods (Guided Backpropagation and Integrated Gradients). The uncertainty estimation methods are organized in individual rows and the explanation methods are aligned along the columns with the columns 1-3 representing the $\mu$, $\sigma$, and the $CV$ heatmaps for IG and the columns 4-6 representing the same for GBP.
5.3 Regression Results
@@
Figure 6: The explanations for the regression task with the mean and the standard deviation explanation (rows 1-2) and the coefficient of variation explanations (rows 3-4). The rows 1 and 3 represent the outputs of GBP and the rows 2 and 4 represent LIME. From left to right, the columns represent as follows: (i) Deep Ensemble (ii) MC-Dropout (iii) MC-DropConnect (iv) Flipout. The input features mapping in the plots is as follows: [1: longitude, 2: latitude, 3: housing median age, 4: total rooms, 5: total bedrooms, 6: population, 7: households, 8: median income].
It can be observed that the mean explanations generated by LIME generally have a lower activation as compared to GBP. This indicates less agreement amongst the constituent explanations and in turn higher uncertainty in the explanation generated by LIME.
Furthermore, we find that for LIME, the value of the coefficient of variation is slightly higher for several features, especially in the case of Deep Ensemble and MC-DropConnect. This implies that these explanations are highly uncertain and should be used with caution. This can be attributed to disagreement among the constituent explanations for that particular feature. The same is observed in the case of GBP for a couple of features, (to a smaller degree as compared to LIME), especially for the case of Deep Ensembles.
@@ 6 Conclusions and Future Work
Both uncertainty estimation and saliency explanations have broad impact on society, as machine learning models are increasingly being used in real-world settings, where humans could be harmed, and uncertainty estimation produces better calibrated output uncertainties to detect when a model makes incorrect or ambiguous predictions. Similarly for saliency explanations, when explanation uncertainty is produced by using an uncertainty estimation method, this would allow human users to obtain additional information in order to decide if the explanation should be trusted or not.
@@
In both cases, output uncertainty and saliency explanations can be misleading, and always additional validation is necessary when used by human users. This paper did not make a user study to evaluate the quality of explanations and uncertainty from a human perspective, we leave this for future work.
Appendix A Additional Model Training Details
The specifications for the model training have been provided in Tables 2 to Table 4.
Hyperparams. CIFAR-10 FER+ CHD
optimizer adam adam RMSProp
loss cross-entropy cross-entropy MSE
metric accuracy accuracy MAE
learning rate 0.001 0.001 0.001
train samples 50000 28559 12750
validation samples 9500 3579 4250
test samples 500 3573 4250
train batch size 2048 2048 4096
validation batch size 2048 2048 4096
Table 2: Hyperparameter used to train the miniVGG16 on CIFAR-10, FER+ and California Housing Dataset (CHD). MSE: Mean Squared Error, MAE: Mean Absolute Error.
Augmentation CIFAR-10 FER+
width shift range 0.1 0.1
height shift range 0.1 0.1
shear range 0.1 -
zoom range 0.1 0.1
horizontal flip True True
rescale 1/255 1/255
rotation range - 30
Table 3: Data augmentation applied to CIFAR-10 and FER+.
Hyperparameter Search Space
# of units in a layer {1, 2, 4, 6, 8}
# of layers {1, 2, 3, 4, 5}
# of epochs {10, 50, 100, 150}
optimizer {‘SGD’, ‘RMSProp’, ‘Adam’}
Table 4: Search space used for the hyperparameter tuning conducted for the MLP.
Figure 7: Parallel plot depicting the results of hyperparameter search conducted for creating an MLP. The optimum values of the hyperparameters are provided in the table below the plot.
Figure 8: Scatter plot depicting the results of hyperparameter search conducted for creating an MLP. The optimum values of the hyperparameters are provided in the table below the plot.
Appendix B Additional Insertion/Deletion Curves
The Area Under Curve (AUCs) for the pixel insertion and deletion metrics for CIFAR-10 and FER+ are provided in Table 5. Additionally, the class specific pixel insertion and deletion curves for FER+ are shown in Figure 10.
Ensemble MC-Dropout MC-DropConnect Flipout
$\mu$-GBP
Deletion
Insertion Refer to caption
Deletion Refer to caption
Insertion Refer to caption
Deletion Refer to caption
Insertion Refer to caption
Deletion Refer to caption
Insertion Refer to caption
$\sigma$-GBP
$\mu$-IG
$\sigma$-IG
Figure 9: Pixel flipping curves for Deep Ensemble (columns 1-2), MC-Dropout (columns 3-4), MC-DropConnect (columns 5-6), and Flipout (columns 7-8) for CIFAR-10. The columns 1, 3, 5, and 7 depict the deletion curves whereas the columns 2, 4, 6, and 8 depict the insertion curves. From top to bottom, the rows show the following heatmaps and the explanation methods they correspond to: (i) $\mu$-GBP (ii) $\sigma$-GBP (iii) $\mu$-IG (iv) $\sigma$-IG.
Ensemble MC-Dropout MC-DropConnect Flipout
$\mu$-GBP
Deletion
Insertion Refer to caption
Deletion Refer to caption
Insertion Refer to caption
Deletion Refer to caption
Insertion Refer to caption
Deletion Refer to caption
Insertion Refer to caption
$\sigma$-GBP
$\mu$-IG
$\sigma$-IG
Figure 10: Pixel flipping curves for Deep Ensemble (columns 1-2), MC-Dropout (columns 3-4), MC-DropConnect (columns 5-6), and Flipout (columns 7-8) for FER+. The columns 1, 3, 5, and 7 depict the deletion curves whereas the columns 2, 4, 6, and 8 depict the insertion curves. From top to bottom, the rows show the following heatmaps and the explanation methods they correspond to: (i) $\mu$-GBP (ii) $\sigma$-GBP (iii) $\mu$-IG (iv) $\sigma$-IG.
CIFAR10 FER+
truck (+ -) frog (+ -) bird (+ -) cat (+ -) deer (+ -) plane (+ -) ship (+ -) auto (+ -) dog (+ -) horse (+ -) contempt (+ -) sadness (+ -) neutral (+ -) surprise (+ -) anger (+ -) happy (+ -) fear (+ -) disgust (+ -)
$\mu$-GBP-Ensemble 0.80 0.62 0.86 0.65 0.48 0.52 0.40 0.47 0.47 0.52 0.34 0.69 0.61 0.53 0.59 0.57 0.30 0.41 0.52 0.56 0.24 0.35 0.25 0.29 0.28 0.36 0.33 0.35 0.31 0.32 0.40 0.44 0.27 0.30 0.29 0.25
$\sigma$-GBP-Ensemble 0.78 0.59 0.87 0.69 0.47 0.51 0.40 0.45 0.45 0.50 0.34 0.66 0.62 0.51 0.59 0.56 0.31 0.39 0.51 0.54 0.25 0.34 0.25 0.28 0.28 0.35 0.33 0.34 0.32 0.32 0.40 0.43 0.27 0.29 0.31 0.24
$\mu$-IG-Ensemble 0.79 0.72 0.86 0.73 0.48 0.53 0.40 0.47 0.46 0.56 0.43 0.62 0.58 0.62 0.57 0.73 0.32 0.43 0.58 0.65 0.31 0.35 0.27 0.30 0.34 0.39 0.38 0.42 0.34 0.35 0.43 0.40 0.32 0.37 0.35 0.33
$\sigma$-IG-Ensemble 0.78 0.66 0.86 0.77 0.48 0.51 0.39 0.45 0.48 0.54 0.40 0.54 0.56 0.59 0.57 0.68 0.32 0.41 0.58 0.56 0.33 0.35 0.27 0.30 0.33 0.38 0.38 0.41 0.34 0.35 0.43 0.41 0.31 0.36 0.35 0.33
$\mu$-GBP-Dropout 0.69 0.29 0.86 0.39 0.56 0.39 0.45 0.25 0.53 0.29 0.37 0.45 0.65 0.33 0.68 0.38 0.30 0.19 0.42 0.24 0.46 0.38 0.34 0.37 0.48 0.48 0.40 0.21 0.46 0.33 0.38 0.22 0.28 0.11 0.39 0.24
$\sigma$-GBP-Dropout 0.70 0.29 0.85 0.40 0.53 0.40 0.44 0.27 0.52 0.32 0.38 0.41 0.62 0.34 0.66 0.38 0.28 0.19 0.43 0.23 0.45 0.36 0.37 0.33 0.48 0.43 0.39 0.21 0.47 0.31 0.41 0.22 0.31 0.18 0.33 0.21
$\mu$-IG-Dropout 0.75 0.38 0.82 0.48 0.54 0.39 0.43 0.28 0.61 0.35 0.47 0.38 0.56 0.40 0.69 0.49 0.39 0.19 0.56 0.34 0.46 0.41 0.36 0.32 0.49 0.43 0.40 0.26 0.46 0.36 0.45 0.24 0.30 0.16 0.45 0.31
$\sigma$-IG-Dropout 0.73 0.36 0.83 0.50 0.59 0.37 0.39 0.27 0.60 0.36 0.47 0.32 0.57 0.41 0.66 0.47 0.36 0.17 0.51 0.31 0.47 0.39 0.37 0.31 0.48 0.42 0.44 0.28 0.44 0.36 0.45 0.25 0.31 0.20 0.44 0.29
$\mu$-GBP-DropConnect 0.71 0.29 0.85 0.36 0.47 0.29 0.50 0.29 0.64 0.25 0.36 0.51 0.69 0.25 0.70 0.29 0.41 0.33 0.52 0.24 0.21 0.29 0.22 0.25 0.23 0.27 0.21 0.23 0.21 0.25 0.27 0.34 0.16 0.19 0.16 0.23
$\sigma$-GBP-DropConnect 0.69 0.30 0.85 0.42 0.45 0.29 0.45 0.30 0.63 0.29 0.33 0.44 0.69 0.25 0.67 0.28 0.37 0.30 0.49 0.23 0.19 0.28 0.23 0.25 0.23 0.27 0.21 0.25 0.23 0.23 0.27 0.36 0.16 0.18 0.16 0.21
$\mu$-IG-DropConnect 0.72 0.38 0.85 0.42 0.50 0.25 0.46 0.32 0.62 0.33 0.46 0.41 0.63 0.33 0.64 0.49 0.43 0.37 0.63 0.35 0.28 0.22 0.30 0.24 0.31 0.26 0.32 0.24 0.26 0.23 0.39 0.28 0.27 0.24 0.31 0.20
$\sigma$-IG-DropConnect 0.72 0.34 0.83 0.46 0.43 0.25 0.38 0.32 0.65 0.36 0.45 0.30 0.62 0.31 0.65 0.47 0.42 0.31 0.60 0.33 0.29 0.25 0.29 0.23 0.30 0.26 0.34 0.24 0.26 0.22 0.37 0.28 0.28 0.26 0.26 0.21
$\mu$-GBP-Flipout 0.70 0.27 0.92 0.39 0.59 0.38 0.51 0.33 0.51 0.22 0.45 0.39 0.59 0.20 0.73 0.35 0.31 0.23 0.55 0.25 0.27 0.32 0.31 0.29 0.32 0.33 0.46 0.27 0.27 0.25 0.32 0.24 0.39 0.28 0.28 0.19
$\sigma$-GBP-Flipout 0.61 0.26 0.92 0.47 0.57 0.40 0.44 0.34 0.51 0.23 0.41 0.32 0.59 0.20 0.69 0.33 0.29 0.22 0.54 0.22 0.30 0.31 0.30 0.30 0.32 0.35 0.45 0.27 0.26 0.26 0.32 0.25 0.36 0.24 0.29 0.21
$\mu$-IG-Flipout 0.70 0.36 0.89 0.49 0.61 0.42 0.48 0.36 0.58 0.29 0.48 0.26 0.52 0.22 0.67 0.44 0.41 0.21 0.64 0.29 0.30 0.31 0.35 0.33 0.40 0.36 0.48 0.33 0.30 0.26 0.44 0.22 0.39 0.33 0.29 0.32
$\sigma$-IG-Flipout 0.69 0.35 0.88 0.52 0.60 0.42 0.43 0.40 0.56 0.29 0.49 0.24 0.56 0.20 0.66 0.44 0.38 0.22 0.64 0.29 0.38 0.32 0.35 0.36 0.40 0.39 0.48 0.35 0.29 0.29 0.47 0.23 0.39 0.32 0.31 0.32
Table 5: Class-wise AUCs for the pixel flipping plots for CIFAR-10 (columns 1-10) and FER+ (columns 11-18). The + and the - denote the pixel insertion and pixel deletion metrics respectively. The combination of the uncertainty estimation methods (Deep Ensemble, MC-Dropout, MC-DropConnect and Flipout, explanation methods (Guided BackPropagation (GBP) and Integrated Gradients (IG)), and type of explanation heatmap (mean ($\mu$) and standard deviation ($\sigma$)) are listed along the rows.
Appendix C Additional Example Explanations
Additional examples of explanation uncertainty (similar to Figures 4 and 5) have been provided in Figures 11 - 16 for both the CIFAR10 and FER+ datasets. It should be noted that the heatmaps in the additional examples are not normalized, unlike the figures in the main paper which have proper normalization.
Input Refer to caption
$E_{\mu}$ IG $E_{\sigma}$ IG CV IG $E_{\mu}$ GBP $E_{\sigma}$ GBP CV GBP
Ensemble Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-D Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-DC Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Example explanations of an airplane from CIFAR-10.
Input Refer to caption
$E_{\mu}$ IG $E_{\sigma}$ IG CV IG $E_{\mu}$ GBP $E_{\sigma}$ GBP CV GBP
Ensemble Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-D Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-DC Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: Example explanations of a deer from CIFAR-10.
Input Refer to caption
$E_{\mu}$ IG $E_{\sigma}$ IG CV IG $E_{\mu}$ GBP $E_{\sigma}$ GBP CV GBP
Ensemble Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-D Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-DC Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: Example explanations of a dog taken from CIFAR-10.
Input Refer to caption
$E_{\mu}$ IG $E_{\sigma}$ IG CV IG $E_{\mu}$ GBP $E_{\sigma}$ GBP CV GBP
Ensemble Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-D Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-DC Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 14: Example explanations of a happy person from FER+.
Input Refer to caption
$E_{\mu}$ IG $E_{\sigma}$ IG CV IG $E_{\mu}$ GBP $E_{\sigma}$ GBP CV GBP
Ensemble Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-D Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-DC Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 15: Example explanations of an angry person from FER+.
Input Refer to caption
$E_{\mu}$ IG $E_{\sigma}$ IG CV IG $E_{\mu}$ GBP $E_{\sigma}$ GBP CV GBP
Ensemble Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-D Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-DC Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 16: Example explanations of a neutral person from FER+.